OpenStudy (anonymous):

Calculate the area of the segment of the curve y^2=4x cut off by the line y=x

OpenStudy (amistre64):

it all in the 1st quadrant so we can equate this to: y = 2 sqrt(x) and y = x right?

OpenStudy (amistre64):

or we can view it from the y axis as: (y^2)/4 = x and y = x

OpenStudy (amistre64):

we need an interval; so use system of equation techniques to determine the values

OpenStudy (amistre64):

(1/4) y^2 - y = 0 y (1/4 y - 1) = 0; y = 0 and y = 4

OpenStudy (amistre64):

given this interval we simply calculate the are under one curve and subtract the given area of the other

OpenStudy (amistre64):

\[\int_{0}^{4}(\frac{1}{4}y^2)dy-\int_{0}^{4}y\ dy\]

OpenStudy (amistre64):

either way you go dx or dy you get 8/3 i believe

OpenStudy (amistre64):

if at any point this makes no sense, just ask